OpenStudy (anonymous):

what expression is equivalent to 3(5 + 6)?

OpenStudy (anonymous):

Do you have choices?

OpenStudy (anonymous):

8 + 9 8 + 6 15 + 18 15 + 6

OpenStudy (anonymous):

Do you know what 5 + 6 is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

then times it by 3

OpenStudy (anonymous):

And what's 3 x 11?

OpenStudy (anonymous):

33

OpenStudy (anonymous):

Good! so find an equation in your choices that equals 11.

OpenStudy (anonymous):

cool thx

OpenStudy (anonymous):

np

OpenStudy (anonymous):

find a equation that equals 33 not 11

OpenStudy (anonymous):

so the correct answer is C 15+18

OpenStudy (anonymous):

Oops, I meant find an answer that equals 33.

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

i look and was like what, that makes no sense so i actually read the question

OpenStudy (anonymous):

8 + 9 = 17 8 + 6 = 14 15 + 18 = 33 15 + 6 = 21 So what is the answer? @Phoenix2003

OpenStudy (anonymous):

C 15 + 18

OpenStudy (anonymous):

Yes. Oh, and you can give people medals by clicking best response next to their comment. You should do it after someone helps you. :)

OpenStudy (anonymous):

i actually got the answer XD

OpenStudy (anonymous):

Thx for letting me know and if you can i have another question going on